# New bounds for proper $h$-conflict-free colourings

### v1

@@ 4. Bounding $\chi_{\rm pcf}^{h}$ for graphs of constrained minimum degree
For a vertex $v\in V_{\leq d}$, the outcome of $B_{{\bm{\sigma}}}(v)$ is entirely determined by the colours assigned to vertices in $N(v)$. So if we fix the realisation of ${\bm{\sigma}}$ outside of $N(v)$, we in particular fix the outcomes of all events $B_{\bm{\sigma}}(u)$ such that $N(v)\cap N[u]=\emptyset$. Hence we let $\Gamma(v)\coloneqq N[N(v)]$, of size $|\Gamma(v)|\leq d\Delta$. We let $\Sigma_{0}$ be the set of possible realisations of ${{\bm{\sigma}}}_{\left|V(G)\setminus N(v)\right.}$. With that definition of $\Gamma(v)$, for every $Z\subset V_{\leq d}\setminus\Gamma(v)$, we infer that
$p\coloneqq\mathbb{P}\left[B_{\bm{\sigma}}(v)\;\middle|\;\bigcap_{u\in Z}\overline{B_{\bm{\sigma}}(u)}\right]\leq\sup_{\sigma_{0}\in\Sigma_{0}}\mathbb{P}\left[B_{{\bm{\sigma}}}(v)\;\middle|\;{{\bm{\sigma}}}_{\left|V(G)\setminus N(v)\right.}=\sigma_{0}\right]\leq\Delta^{-2.6}.$
Since $\Delta\geq\delta\geq 16h\geq 320\log\Delta$, we must have $\Delta>2500$. So we have $ep(d\Delta)\leq e\Delta^{-0.6}<1$. We may therefore apply Lemma 17 to conclude that, with non-zero probability, no event $B_{{\bm{\sigma}}}(v)$ occurs. ∎
A direct application of Lemma 25 with $d\coloneqq\Delta$ yields the following result as a corollary.
@@
[…]
@@
We execute the following procedure, which is identical to the one in the proof of Lemma 19 except for the fact that we no longer ensure that witnesses in $\sigma_{0}$ remain witnesses in the final colouring.
@@
For $i\in[m]$, we respectively let $\mathbf{A}_{i}$, $\mathbf{S}_{i}$, $\mathbf{C}_{i}$, and $\mathbf{i_{0}}$ be the random values of the variables $A_{i}$, $S_{i}$, $C_{i}$, and $i_{0}$ at the end of the procedure. Let ${\bm{\sigma}}$ be the colouring obtained from $\sigma_{0}$ after recolouring $\mathbf{C}_{i}$ with the colour $k+i$ for each $i\in\{\mathbf{i_{0}},\dots,m\}$. As explained in the proof of Lemma 19, the colouring ${\bm{\sigma}}$ is deterministically proper. For $v\in V_{>d}$, let $B_{v}$ be the (bad) event that more than $\deg(v)/2$ neighbours of $v$ belong in $\mathbf{A}_{1}\cup\dots\cup\mathbf{A}_{m}$.
@@
Let $\mathbf{Z}_{v}\coloneqq\sum_{i=1}^{m}|\mathbf{A}_{i}\cap N(v)|\sim B(m\deg(v),p)$. By Lemma 14, we obtain
@@
where we have used Equation 2 and the assumption $\deg(v)\geq\delta\geq 20\log\Delta$. ∎
@@
Let $\mathbf{Y}_{v}\coloneqq\sum_{i=1}^{m}|\mathcal{W}(v)\cap A_{i}|\sim B(mh_{0},p)$. By Lemma 14, we obtain
@@
where we have used Equation 2 and the assumption $h_{0}\geq 20\log\Delta$. ∎
For $v\in V_{>d}$, let $\mathbf{X}_{v}^{i}$ be the indicator variable for the event “$|N(v)\cap\mathbf{C}_{i}|=1$”. The following fact is easily deduced from the disjointness of the sets $\mathbf{C}_{\mathbf{i_{0}}},\dots,\mathbf{C}_{m}$.
@@
Let $\mathbf{X}_{v}\coloneqq\sum_{i=1}^{m}\mathbf{X}_{v}^{i}$. If $\mathbf{i_{0}}=1$ and $\mathbf{X}_{v}\geq\eta$, then $v$ has at least $\eta$ witnesses in ${\bm{\sigma}}$, as desired. Let $E_{v}$ be the (bad) event that $\mathbf{X}_{v}<\eta$.
@@
Consider $i\in[m]$. Let us fix the realisation $\mathbf{A}_{1},\dots,\mathbf{A}_{i-1}$ and the resulting value of $\mathbf{C}_{1},\dots,\mathbf{C}_{i-1}$. Let $C^{\prime}$ be the resulting value of the term “$\bigcup_{j=i_{0}}^{i-1}C_{j}$” at the start of round $i$. The distribution of $\mathbf{A}_{i}$ is unaffected by this conditioning, i.e. every vertex of $v$ is still present in $\mathbf{A}_{i}$ independently with probability $p$. Let $U\coloneqq N(v)\setminus C^{\prime}$ and let $t\coloneqq|U|$. Because of the fail-safe, $C^{\prime}$ contains at most $\deg(v)/2$ neighbours of $v$, therefore $t\geq\deg(v)-\frac{\deg(v)}{2}\geq\frac{d}{2}$.
The probability that $N(v)\cap\mathbf{A}_{i}=\{u\}$ with $u\in U$ is equal to $tp(1-p)^{\deg(v)-1}\geq\frac{dp}{2}(1-p)^{\Delta-1}$. Suppose such a vertex $u\in U$ exists, i.e. we are conditioning on the event “$N(v)\cap\mathbf{A}_{i}=\{u\}$ for some $u\in U$”. By independence of the random choices, the distribution of $\mathbf{A}_{i}$ outside of $N(v)$ remains unaffected, i.e. every vertex $x\in V\setminus N(v)$ is present in $\mathbf{A}_{i}$ independently with probability $p$. We now determine a lower bound on the probability that $u\in\mathbf{C}_{i}$, which is equivalent to $u\notin\mathbf{S}_{i}$ since $u\in\mathbf{A}_{i}$. By definition of $U$, $u\notin C^{\prime}$, so $u\notin\mathbf{S}_{i}$ if and only $N(u)\cap\mathbf{A}_{i}=\emptyset$: we already know that $N(v)\cap\mathbf{A}_{i}=\{u\}$, therefore this happens with probability $(1-p)^{|N(u)\setminus N(v)|}\geq(1-p)^{\Delta}$. Therefore,
@@
where we have used the fact that $(1-\frac{1}{x})^{x-1}>\frac{1}{e}$ for all $x\geq 1$. Let $q\coloneqq\frac{dp}{2e}$. We have shown that $\mathbb{P}\left[\mathbf{X}_{v}^{i}=1\right]>q$ regardless of the realisation of $\mathbf{A}_{1},\dots,\mathbf{A}_{i-1}$, and so in particular,
@@
By Lemma 15, we obtain $\mathbb{P}\left[\sum\mathbf{X}_{v}^{i}<\eta\right]\leq\mathbb{P}\left[B(m,q)<\eta\right]$. We have defined $m$ such that $mq\geq 2\eta$, and since $\eta\geq 20\log\Delta$ we have $\frac{mq}{8}\geq 5\log\Delta$; by Lemma 13(i), we conclude that $\mathbb{P}\left[E_{v}\right]\leq\mathbb{P}\left[B(m,q)<\eta\right]\leq\Delta^{-5}$. ∎
The bad events $B_{v}$ and $E_{v}$ are associated to vertices of $V_{>d}$, while the bad events $E^{-}_{v}$ are associated to vertices of $V_{\leq d}$. Each of these events are determined by the realisation of $(\mathbf{A}_{1},\ldots,\mathbf{A}_{m})$ within the distance $2$ neighbourhood of their corresponding vertices, therefore these bad events are independent from all but at most $2\Delta^{4}$ other such bad events. By Lemma 16, with nonzero probability, none of these bad events occur. In that case, $\mathbf{i_{0}}=1$ and $\mathbf{X}_{v}\geq\eta$ for every $v\in V_{>d}$, while $|\mathcal{W}(v)\setminus(\mathbf{A}_{1}\cup\dots\cup\mathbf{A}_{m})|\geq h_{0}/2$ for every $v\in V_{\leq d}$. By 29 and 31, we conclude that the colouring ${\bm{\sigma}}$ is $\langle V_{\leq d}\;?\;h_{0}/2:\eta\rangle$-conflict-free.
@@
Let $G$ be a graph of maximum degree $\Delta\geq 40000$ and minimum degree $\delta\geq 3000\log\Delta$. Let $h$ be an integer such that $10020\log\Delta\leq h\leq\delta/$ 100. If $\delta\leq\sqrt{h\Delta}$, then
@@
[…]
@@
In the statement of 33, we did not consider the case $\delta>\sqrt{h\Delta}$. In that case, 9 provides the upper bound $\chi_{\rm pcf}^{h}(G)\leq\Delta+30\frac{h\Delta}{\delta}$; observe that the second-order term is $\mathcal{O}(\sqrt{h\Delta})$, and approaches $\mathcal{O}(h)$ as $\delta\rightarrow\Delta$.
@@ 5. Discussion
As discussed in the introduction, the bound $\Delta+\mathcal{O}(\sqrt{h\Delta})$ is unlikely to be tight. When $h=1$, we have a bound of the form $\Delta+O(\log\Delta)$, while when $h=2$ we have no better than $\Delta+\mathcal{O}(\sqrt{\Delta\log\Delta})$ with minimum degree $\Omega(\log\Delta)$. We expect the transition to be smoother between $h=1$ and larger values for $h$, as is suggested by 11. At the other end of the spectrum, when $h$ gets close to $\delta$, it has been established in [7] that a linear upper bound in terms of $\Delta$ is no longer possible, at least for regular graphs. This was formulated for $\chi_{o}^{h}(G)$, which we recall is a lower bound on $\chi_{\rm pcf}^{h}(G)$.
@@
Let $G$ be a $\Delta$-regular graph, let $\left\lceil 2(\ln\Delta+\ln\ln\Delta=3\right\rceil\leq t\leq\Delta$ be a given integer, and let $h\coloneqq\Delta+1-t$. Then
@@
Let $h\coloneqq\Delta-t$, and let $\sigma$ be any $h$-odd colouring of $G$. Let $v\in V(G)$; we say that $c\in\sigma(N(v))$ is an odd colour of $v$ in $\sigma$ if it has an odd number of occurrences in $N(v)$. Let $x\geq 0$ be the number of odd colours of $v$ that have $>1$ occurrences in $N(v)$. There can be at most $\Delta-3x$ other (odd) colours in $N(v)$, hence $h\leq\Delta-2x$. We conclude that $x\leq\frac{\Delta-h}{2}=t/2$, and so $v$ has at least $\Delta-3t/2$ solitary colours in $\sigma$. This holds for every $v\in V(G)$, so $\sigma$ is $(\Delta-3t/2)$-conflict-free. ∎

### v2

@@ 4. Bounding $\chi_{\rm pcf}^{h}$ for graphs of constrained minimum degree
For a vertex $v\in V_{\leq d}$, the outcome of $B_{{\bm{\sigma}}}(v)$ is entirely determined by the colours assigned to vertices in $N(v)$. So if we fix the realisation of $\bm{\sigma}$} outside of $N(v)$, we in particular fix the outcomes of all events $B_{\bm{\sigma}}(u)$ such that $N(v)\cap N[u]=\emptyset$. Hence we let $\Gamma(v)\coloneqq N[N(v)]\setminus\{v\}$, of size $|\Gamma(v)|\leq d\Delta$. We let $\Sigma_{0}$ be the set of possible realisations of ${{\bm{\sigma}}}_{\left|V(G)\setminus N(v)\right.}$. With that definition of $\Gamma(v)$, for every $Z\subset V_{\leq d}\setminus\Gamma(v)$, we infer that
$\mathbb{P}\left[B_{\bm{\sigma}}(v)\;\middle|\;\bigcap_{u\in Z}\overline{B_{\bm{\sigma}}(u)}\right]\leq\sup_{\sigma_{0}\in\Sigma_{0}}\mathbb{P}\left[B_{{\bm{\sigma}}}(v)\;\middle|\;{{\bm{\sigma}}}_{\left|V(G)\setminus N(v)\right.}=\sigma_{0}\right]\leq\Delta^{-2.6}\eqqcolon p.$
Since $\Delta\geq\delta\geq 16h\geq 320\log\Delta$, we must have $\Delta>2500$. So we have $4p(d\Delta)\leq 4\Delta^{-0.6}<1$. We may therefore apply Lemma˜17 to conclude that, with non-zero probability, no event $B_{{\bm{\sigma}}}(v)$ occurs. ∎
A direct application of Lemma˜25 with $d\coloneqq\Delta$ yields the following result as a corollary.
@@
There remains to show how to transform a $\langle V_{\leq d}\;?\;2h:0\rangle$-conflict-free colouring into a conflict-free colouring through a random recolouring process. The main idea follows from Lemma˜19 with a better adaptation to the present situation.
@@
We execute the following procedure, which is identical to the one in the proof of Lemma˜19 except for the fact that we no longer ensure that witnesses in $\sigma_{0}$ remain witnesses in the final colouring.
@@
For $i\in[m]$, we respectively let $\mathbf{A}_{i}$, $\mathbf{S}_{i}$, $\mathbf{C}_{i}$, and $\mathbf{i_{0}}$ be the random values of the variables $A_{i}$, $S_{i}$, $C_{i}$, and $i_{0}$ at the end of the procedure. Let $\bm{\sigma}$} be the colouring obtained from $\sigma_{0}$ after recolouring $\mathbf{C}_{i}$ with the colour $k+i$ for each $i\in\{\mathbf{i_{0}},\dots,m\}$. As explained in the proof of Lemma˜19, the colouring $\bm{\sigma}$} is deterministically proper. For $v\in V_{>d}$, let $B_{v}$ be the (bad) event that more than $\deg(v)/2$ neighbours of $v$ belong in $\mathbf{A}_{1}\cup\dots\cup\mathbf{A}_{m}$.
@@
Let $\mathbf{Z}_{v}\coloneqq\sum_{i=1}^{m}|\mathbf{A}_{i}\cap N(v)|\sim B(m\deg(v),p)$. By Lemma˜14, we obtain
@@
where we have used Equation˜2 and the assumption $\deg(v)\geq\delta\geq 20\log\Delta$. ∎
@@
Let $\mathbf{Y}_{v}\coloneqq\sum_{i=1}^{m}|\mathcal{W}(v)\cap A_{i}|\sim B(mh_{0},p)$. By Lemma˜14, we obtain
@@
where we have used Equation˜2 and the assumption $h_{0}\geq 20\log\Delta$. ∎
For $v\in V_{>d}$, let $\mathbf{X}_{v}^{i}$ be the indicator variable for the event $|N(v)\cap\mathbf{C}_{i}|=1$”. The following fact is easily deduced from the disjointness of the sets $\mathbf{C}_{\mathbf{i_{0}}},\dots,\mathbf{C}_{m}$.
@@
For every $v\in V_{>d}$, let $\mathbf{X}_{v}\coloneqq\sum_{i=1}^{m}\mathbf{X}_{v}^{i}$. If $\mathbf{i_{0}}=1$ and $\mathbf{X}_{v}\geq\eta$, then $v$ has at least $\eta$ witnesses in $\bm{\sigma}$}, as desired. Let $E_{v}$ be the (bad) event that $\mathbf{X}_{v}<\eta$.
@@
Consider $i\in[m]$. Let us fix the realisation $\mathbf{A}_{1},\dots,\mathbf{A}_{i-1}$ and the resulting value of $\mathbf{C}_{1},\dots,\mathbf{C}_{i-1}$. Let $C^{\prime}$ be the resulting value of the term $\bigcup_{j=i_{0}}^{i-1}C_{j}$” at the start of round $i$. The distribution of $\mathbf{A}_{i}$ is unaffected by this conditioning, i.e. every vertex of $v$ is still present in $\mathbf{A}_{i}$ independently with probability $p$. Let $U\coloneqq N(v)\setminus C^{\prime}$ and let $t\coloneqq|U|$. Because of the fail-safe, $C^{\prime}$ contains at most $\deg(v)/2$ neighbours of $v$, therefore $t\geq\deg(v)-\frac{\deg(v)}{2}\geq\frac{d}{2}$.
The probability that $N(v)\cap\mathbf{A}_{i}=\{u\}$ with $u\in U$ is equal to $tp(1-p)^{\deg(v)-1}\geq\frac{dp}{2}(1-p)^{\Delta-1}$. Suppose such a vertex $u\in U$ exists, i.e. we are conditioning on the event $N(v)\cap\mathbf{A}_{i}=\{u\}$ for some $u\in U$”. By independence of the random choices, the distribution of $\mathbf{A}_{i}$ outside of $N(v)$ remains unaffected, i.e. every vertex $x\in V\setminus N(v)$ is present in $\mathbf{A}_{i}$ independently with probability $p$. We now determine a lower bound on the probability that $u\in\mathbf{C}_{i}$, which is equivalent to $u\notin\mathbf{S}_{i}$ since $u\in\mathbf{A}_{i}$. By definition of $U$, $u\notin C^{\prime}$, so $u\notin\mathbf{S}_{i}$ if and only $N(u)\cap\mathbf{A}_{i}=\emptyset$. We already know that $N(v)\cap\mathbf{A}_{i}=\{u\}$, therefore this happens with probability $(1-p)^{|N(u)\setminus N(v)|}\geq(1-p)^{\Delta}$. Therefore,
@@
where we have used the fact that $(1-\frac{1}{x})^{x-1}>\frac{1}{e}$ for all $x>1$. Let $q\coloneqq\frac{dp}{2e}$. We have shown that $\mathbb{P}\left[\mathbf{X}_{v}^{i}=1\right]>q$ regardless of the realisation of $\mathbf{A}_{1},\dots,\mathbf{A}_{i-1}$, and so in particular,
@@
By Lemma˜15, we obtain $\mathbb{P}\left[\sum\mathbf{X}_{v}^{i}<\eta\right]\leq\mathbb{P}\left[B(m,q)<\eta\right]$. We have defined $m$ such that $mq\geq 2\eta$, and since $\eta\geq 20\log\Delta$ we have $\frac{mq}{8}\geq 5\log\Delta$; by Lemma˜13(i), we conclude that $\mathbb{P}\left[E_{v}\right]\leq\mathbb{P}\left[B(m,q)<\eta\right]\leq\Delta^{-5}$. ∎
The bad events $B_{v}$ and $E_{v}$ are associated to vertices of $V_{>d}$, while the bad events $E^{-}_{v}$ are associated to vertices of $V_{\leq d}$. Each of these events are determined by the realisation of $(\mathbf{A}_{1},\ldots,\mathbf{A}_{m})$ within the distance $2$ neighbourhood of their corresponding vertices, therefore these bad events are independent from all but at most $2\Delta^{4}$ other such bad events. By Lemma˜16, with nonzero probability, none of these bad events occur. In that case, $\mathbf{i_{0}}=1$ and $\mathbf{X}_{v}\geq\eta$ for every $v\in V_{>d}$, while $|\mathcal{W}(v)\setminus(\mathbf{A}_{1}\cup\dots\cup\mathbf{A}_{m})|\geq h_{0}/2$ for every $v\in V_{\leq d}$. By ˜29 and ˜31, we conclude that the colouring $\bm{\sigma}$} is $\langle V_{\leq d}\;?\;h_{0}/2:\eta\rangle$-conflict-free. ∎
@@
Let $G$ be a graph of maximum degree $\Delta\geq 20000$ and minimum degree $\delta\geq 2000\log\Delta$. Let $h$ be an integer such that $20\log\Delta\leq h\leq\delta/100$. If $\delta\leq\sqrt{h\Delta}$, then
@@
Let $d\coloneqq\left\lceil 3\sqrt{h\Delta}\right\rceil$. By Lemma˜25, there exists a $(\Delta+3d)$-colouring $\sigma_{0}$ that is proper and $\langle V_{\leq d}\;?\;2h:0\rangle$-conflict-free. We apply Lemma˜27 with $k\coloneqq\Delta+3d$, $h_{0}\coloneqq 2h$ and $\eta\coloneqq h$; the condition $\eta\leq d/100$ is verified by the assumptions $h\leq\delta/100$ and $\delta\leq d$. Then there exists an $h$-pcf $k$-colouring of $G$, with $k=\Delta+3d+\left\lceil 8e\frac{\Delta}{d}h\right\rceil\leq\Delta+17\sqrt{h\Delta}$. ∎
@@
In the statement of ˜33, we did not consider the case $\delta>\sqrt{h\Delta}$. In that case, ˜9 provides the upper bound $\chi_{\rm pcf}^{h}(G)\leq\Delta+30\frac{h\Delta}{\delta}$; observe that the second-order term is $\mathcal{O}(\sqrt{h\Delta})$, and approaches $\mathcal{O}(h)$ as $\delta\rightarrow\Delta$.
@@ 5. Discussion
5.1. Tightness of the bounds
As discussed in the introduction, the bound $\Delta+\mathcal{O}(\sqrt{h\Delta})$ is unlikely to be tight. When $h=1$, we have a bound of the form $\Delta+O(\log\Delta)$, while when $h=2$ we have no better than $\Delta+\mathcal{O}(\sqrt{\Delta\log\Delta})$ with minimum degree $\Omega(\log\Delta)$. We expect the transition to be smoother between $h=1$ and larger values for $h$, as is suggested by ˜11. At the other end of the spectrum, when $h$ gets close to $\delta$, it has been established in [7] that a linear upper bound in terms of $\Delta$ is no longer possible, at least for regular graphs. This was formulated for $\chi_{o}^{h}(G)$, which we recall is a lower bound on $\chi_{\rm pcf}^{h}(G)$.
@@ 5.1. Tightness of the bounds
Let $G$ be a $\Delta$-regular graph, let $\left\lceil 2(\ln\Delta+\ln\ln\Delta+3)\right\rceil\leq t\leq\Delta$ be a given integer, and let $h\coloneqq\Delta+1-t$. Then
@@
Let $h\coloneqq\Delta-t$, and let $\sigma$ be any $h$-odd colouring of $G$. Let $v\in V(G)$; we say that $c\in\sigma(N(v))$ is an odd colour of $v$ in $\sigma$ if it has an odd number of occurrences in $N(v)$. Let $x\geq 0$ be the number of odd colours of $v$ that have >1>1 occurrences in $N(v)$. There can be at most $\Delta-3x$ other (odd) colours in $N(v)$, hence $h\leq\Delta-2x$. We conclude that $x\leq\frac{\Delta-h}{2}=t/2$, and so $v$ has at least $\Delta-3t/2$ solitary colours in $\sigma$. This holds for every $v\in V(G)$, so $\sigma$ is $(\Delta-3t/2)$-conflict-free. ∎
5.2. Algorithmic aspects
We note that the proof of Lemma˜18 is constructive with the description of a greedy algorithm that returns the desired solution. While the statements of Lemma˜19 and Lemma˜27 are purely existential, the only non-constructive part in their respective proofs is the application of Lemma˜16. If one replaces the latter with the following algorithmic version of the Lovász Local Lemma due to Moser and Tardos [13], then both these results can be obtained constructively with a randomised algorithm of expected polynomial time.
Lemma 37 (Algorithmic Lovász Local Lemma).
Let $\mathcal{B}=\{B_{1},\ldots,B_{n}\}$ be a finite set of random (bad) events that are all determined by a finite set of random independent variables $\mathcal{X}$, and all of which happen with probability at most $p$. Suppose that, for every $i\in[n]$, $B_{i}$ is independent from all but at most $d$ other events.
If $epd\leq 1$, then there exists a valuation of the variables in $\mathcal{X}$ such that no bad event in $\mathcal{B}$ happens, and such a valuation can be constructed with a randomised algorithm of expected runtime polynomial in $n$ and $|\mathcal{X}|$.
Indeed, in the proofs of Lemma˜19 and Lemma˜27, the bad events are all determined by the set of independent random variables $\mathcal{X}\coloneqq\{\mathds{1}_{v\in A_{i}}:i\in[m],v\in V(G)\}$. We conclude that one can obtain the colouring promised by ˜5 in expected polynomial time.
In contrast, the proof of Lemma˜25 relies on the lopsided version of LLL (Lemma˜17) which seems to be much harder to prove constructively (see [9] for a partial result in that direction), and more crucially on a much more complex random object, namely a uniformly random proper colouring $\bm{\sigma}$} of the graph $G$. Sampling $\bm{\sigma}$} — or a good enough approximation thereof — in polynomial time is by itself a notoriously hard open problem. If one wishes to make that proof constructive, the most promising strategy in our opinion would be to apply directly the Entropy Compression Method used by Moser and Tardos in their proof of Lemma˜37 on the following random algorithm, which we conjecture should return the desired colouring in expected polynomial time.
Fix an arbitrary ordering on $V(G)$
$k\leftarrow\Delta+3d$
$\sigma\leftarrow$ any proper $k$-colouring of $G$ (constructed greedily)
while there is a vertex $v$ of degree at most $d$ with less than $2h$ witness in $\sigma$ do
 foreach colour class $C$ in $N(v)$ do
    foreach $u\in C\setminus\min C$ do
       $\sigma(u)\leftarrow$ uniformly random colour from $k]\setminus\sigma(N(u))$.
return $\sigma$
Since the algorithmic aspects of our proofs are not the focus of the paper, we leave that for future work.